OpenStudy (shortypants):

For the function ​g(x)=8​, evaluate the following. ​(a)​ g(5​)= ​(b) ​g(−150​)= ​(c) ​g(q​)= ​(d) ​g(2−3t​)=

OpenStudy (magallar23):

The answer is 8. For all of them.

OpenStudy (shortypants):

can you explain why?

OpenStudy (magallar23):

They are asking you what the function value is when you replace the x.

OpenStudy (magallar23):

if you plug in whatever value for x, will it affect the 8?

OpenStudy (shortypants):

well the 2-3t i thought would

OpenStudy (shortypants):

wouldn't it look like 2-3t=8

OpenStudy (magallar23):

It will be f(2-3t) = 8

OpenStudy (shortypants):

which could be solved? right

OpenStudy (magallar23):

You did. You're solving the right hand side. Since nothing changed. It is still 8. The answer is then 8.

OpenStudy (shortypants):

-3t=6

OpenStudy (shortypants):

oh ok

OpenStudy (shortypants):

thanks for your help

OpenStudy (magallar23):

​(a)​ g(5​)= ​(b) ​g(−150​)= ​(c) ​g(q​)= ​(d) ​g(2−3t​)= Everything inside the parentheses is replacing the x. since there is no x in the right side then you're not replacing anything on the right side. X can be anything and it won't change the right hand side.

OpenStudy (shortypants):

thank you that clears it up alot